Solve by completing the square:
x^2 + 2x - 8 = 0
Multiple choice:
a)-4, -2
b)-4, 2
c)4, -2
d)4,2
Thank you for helping!

Answer by ankor@dixie-net.com(22740)

Solve by completing the square:
x^2 + 2x - 8 = 0
:
x^2 + 2x + ___ = 8; added 8 to both sides
:
Find the value that completes the square.
Divide the coefficient of x by 2, and square that value: 2/2 = 1; 1^2 = 1
:
x^2 + 2x + 1 = 8 + 1; we have to add 1 to both sides
:
(x + 1)^2 = 9; Note that (x+1)^2 is same as:(x+1)*(x+1), which = x^2 + 2x + 1)
:
Find the square root of both sides:
x + 1 = +/- Sqrt(9)
:
x = -1 +/- Sqrt(9)
:
x = -1 + 3
x = +2
and
x = -1 - 3
x = -4
:
Can you do this now?